OpenStudy (anonymous):

use limit rule to find derivative of square root of x+2

OpenStudy (anonymous):

\[f(x)= x+4\] ......(1) \[f(x+h)=x+h+4=(x+4)+h=f(x)+h\] ......(from 1) \[\frac{ dy }{ dx }= \lim_{h \rightarrow 0}\frac{ f(x+h) - f(x) }{ h } =\lim_{h \rightarrow 0}\frac{ f(x)+h- f(x) }{ h }=\lim_{h \rightarrow 0}\frac{ h }{ h }=1\]

OpenStudy (anonymous):

I knows how to do it with polynomials but anything that is not polynomials confuses me especially when it has square root

OpenStudy (anonymous):

For example???